Please help with all parts... I am a bit confused and the prof is no help.20.) A newly discovered oxygen binding protein (OBP) has been discovered. The Hill plot shown below was generated based on th
Answer #1

20.

(a) Myoglobin and haemoglobin are two oxygen binding protein which has higher affinity with oxygen.

(b). To calculate the hill Coefficient we must have to use the formula

Log(YO2/1-YO2) = n log po2 - n log p50

So first we have to calculate p50 which is according to this graph,

The constant P50 is the value of pO2 that gives occupancy = 0.5,. P50 = 0.26 for this.

So, n = log(YO2/1 - YO2) + log p50 ÷ log po2

= log(1) + log (0.26) ÷ log(0.26)

= 1

(c). The constant P50 is the value of pO2 that gives occupancy = 0.5. P50 = 0.26 torr for this graph.

(d) . To calculate the percentage saturation we can use the formula.

Y =( po2 / p50 + po2) ×100

. = (10/ 0.26+10) ×100

. = 0.97 × 100

. = 97 %